OpenStudy (anonymous):

If a 65kg woman is driving off a 10 meter platform, what is her speed as she hits the water?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Use:\[v^2=u^2+2as\]

OpenStudy (nowhereman):

Why does this has to do anything with her weight? Also this is a question for the physics section I would say. (hint: earth speed up is \(9.81 m/s^2\))

OpenStudy (anonymous):

\[v^2=u^2+2as\] Where u=initial velocity=0m/s a=g=9.81m/s^2 s=10m

OpenStudy (anonymous):

information not clear

OpenStudy (anonymous):

\[v^2=u^2+2as\]u=0, a=g=9.81ms^-2, s=h (height). So:\[v^2=0^2+2 \times g \times h=2 \times 9.81 \times10=196.2\](Don't worry about units there). So:\[v=\sqrt{196.2}\approx14ms^{-1}\]